[Brahms-l] Shift report 20050212 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 12 2005 - 07:57:48 EST
1. Two stores (0000-0500 and 0710-????), had problems with RHIC inbetween.
2. Runs:
FS@3deg@3450A, MRS@40deg@2255B
13115 - 200k tr.2, 100k tr.3
13116 - 190k tr.2,  90k tr.3
13125 - 277k tr.2, 213k tr.3
Pedestals: 13127, 13128.
13129 - running.
3. We discovered that there is some problems with trigger 6 - see elog for details. There should be expert checking on this problem.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 12 07:58:07 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 12 2005 - 07:58:15 EST